OpenStudy (anonymous):

a plane flying into the wind takes about one hour to fly between Nashville and Lexington .It returns with the wind at its back in 50 min. Find the speed of the plane if there are 320 miles between the cities.

OpenStudy (anonymous):

352 miles per hour

OpenStudy (anonymous):

u got it?

OpenStudy (anonymous):

or i have to elaborate?

OpenStudy (nikvist):

\[v-w=\frac{s}{t_1}\]\[v+w=\frac{s}{t_2}\]\[v=\frac{s}{2}\frac{t_1+t_2}{t_1t_2}=\frac{320}{2}\frac{1+5/6}{5/6}\frac{miles}{h}=352\frac{miles}{h}\]